### v1

# Advection of the image point in probabilistically-reconstructed phase spaces

Abstract
Insufficient reference data is ubiquitous in data-driven computational fluid dynamics, as it is usually too expensive to compute or impossible to observe over long enough times needed for data-driven methods to properly operate out of the sampled range. Ultimately, the lack of data can significantly compromise the fidelity of results computed with data-driven methods or make the data-driven approach inapplicable. In order to challenge this problem, we propose a probabilistic reconstruction method which enriches the hyper-parameterisation approach with ideas underlying the probabilistic-evolutionary approach. We offer to use Advection of the image point (a hyper-parameterisation method) on data sampled from the joint probability distribution of the reference dataset. The idea is to blend together the simplicity of the hyper-parameterisation (HP) method with an extra source of reference data provided through sampling from the joint probability distribution.
The HP method has been tested on the sea surface temperature and surface relative vorticity computed with the global 1/4 and 1/12 resolution NEMO model. Our results show that the HP solution (the solution computed with the HP method) in the probabilistically-reconstructed and reduced (in terms of dimensionality) phase space at 1/4 resolution is more accurate compared to the 1/4-solution computed with NEMO and it is also several orders of magnitude faster to compute than the 1/4-run. The proposed method shows encouraging results for the NEMO model and the potential for the use in other operational ocean and ocean-atmospheric models for both deterministic and probabilistic predictions.
[…]
The importance of data-driven modelling in science and engineering is hard to underestimate. However, its applicability can be significantly limited due to the problem of data-deficiency. This problem can lead to inaccurate, low-fidelity results, or simply to inapplicability of the data-driven approach. The proposed study challenges the data-deficiency problem by offering to sample from the joint probability distribution of reference data to enrich the original reference data with new data from the reference data distribution. The method proposed in this work allows to harvest extra reference data to enable data-driven methods to work with meager datasets on which they would be inapplicable otherwise.
1 Introduction
The data-driven paradigm is gradually gaining momentum in computational fluid dynamics by offering an alternative to the physics-driven approach in situations when there is no physics-based model for the studied phenomenon or it is too computationally-expansive to acquire data from the model (e.g. (Mendez et al., 2023) and references therein). Being a powerful tool providing a wide spectrum of methods (e.g. Schneider et al. (2017); Zhang and Lin (2018); O’Gorman and Dwyer (2018); Zanna and Bolton (2020)) ranging from purely-data to physics-data driven, also know as hybrids, the data-driven approach typically suffers from the lack of reference data that can significantly compromise its fidelity or make it inapplicable. In order to help this problem, we offer a probabilistic reconstruction method which augments the hyper-parameterisation (HP) approach (e.g., Shevchenko and Berloff (2021, 2022a, 2022b, 2023a); Shevchenko and Crisan (2024)) with some ideas underlying the probabilistic-evolutionary approach (Shevchenko and Berloff, 2023b). More specifically, we propose to use Advection of the image point (a purely data-driven HP method working in phase space, also called state space) together with data sampled from the joint probability distribution of reference data so as to probabilistically reconstruct the reference phase space for the HP method and make it operable for short reference data records otherwise falling out of its range of applicability. Although, we consider the use of probabilistically-reconstructed phase spaces within the HP framework, its application is not bound to HP. For example, probabilistically-reconstructed reference datasets can be used within the context of other data-driven methods, as well as for training artificial neural networks, and for reconstruction of missing parts in observational data from different sources (drifters, weather stations, satellites, etc.)
2 The probabilistic reconstruction method: schematic and basics
The idea underlying the probabilistic reconstruction method is to calculate the joint probability distribution (JPD) for a given reference dataset and then sample reference data from this JPD instead of using the reference dataset itself. Hence, the JPD becomes an extra layer between the reference data and a data-driven method thus providing an extra source of reference data from the reference data distribution. It would be instructive to first look into the schematic of the proposed method (Fig. 1) for better understanding of its specifics and then consider the method in more detail.
Figure 1: Schematic of the probabilistic reconstruction method.
(a) Reference data acquisition. The first step is to provide the method with reference data. Depending on the problem to solve, it can be numerical simulations, observations, or both. In our case it is sea surface temperature (SST) computed with a global 1/12 resolution ocean model NEMO and then interpolated onto a 1/4 grid (Fig. 1a). If observations are available then they can also be included in the reference dataset by simply adding them to the reference data record. In order to be consistent in space, observations must be on the same grid as the NEMO data; consistency in time is not required, as the proposed method does not use it.
(b) Reference phase space calculation. The second step is to get the reference phase space for the HP method called ”Advection of the image point” (Shevchenko and Berloff, 2021, 2023a); the reference phase space is shown as an orange blob in Fig. 1b. The latter needs locations (which is SST itself) and directional vectors (SST tendencies), which are calculated from the reference data; different schemes can be used, we use the central finite difference in time. The tendencies are shown as black vectors in Fig. 1b, and the points they are attached to represent SST. The white spots (which we call voids) are regions of the phase space which lack data (usually, due to short reference records or unreliable observations). Typically, these voids is a reason for why data-driven methods cease to work on meager phase spaces.
(c) Joint Probability Distribution calculation. The third step is to calculate the joint probability distribution (JPD) from the reference data in phase space; it is shown as a blue surface in Fig. 1c for illustration purposes; in high-dimensional phase spaces (like the one used for SST computed with NEMO) the JPD is a hyper-surface. It is important to remark that we calculate the JPD only for SST, while SST tendencies are computed in a different way explained below. The JPD for the reference SST is calculated globally with the histogram method, i.e. the whole reference dataset contributes to the histogram calculation. As the proposed method is developed to work in high-dimensional phase spaces, computing a multidimensional JPD and keeping it in memory for further sampling is an unaffordable option. Instead, we calculate a coordinate-wise probability density function (PDF). It gives us access to all necessary information for sampling at virtually no cost; we refer the reader to (Shevchenko and Berloff, 2023b) for a detail discussion on the JPD calculation. The only difference in the JPD calculation in this study is that we compute the JPD globally (i.e. for all reference data), while in (Shevchenko and Berloff, 2023b) it is computed locally in a given neighborhood.
(d) Sampling from JPD and calculation of probabilistically-reconstructed phase space. In the fourth step we sample SST from the JPD and then calculate SST tendencies by averaging reference SST tendencies over the neighborhood of sampled SST (Fig. 2). The localized calculation of tendencies allows to decrease the dimensionality of the JPD two times (as the JPD dimension for both SST and its tendencies is two times larger compared to the JPD only for SST) and thus make its calculation more efficient without compromising on accuracy. It is worth mentioning that tendencies can also be computed in a probabilistic way from a local JPD which is constructed based on the information from the neighborhood of the new tendency (red contour in Fig. 2), but this method is slower than the one used in this study.
Figure 2: Shown is a schematic representation of the probabilistically-reconstructed phase space with the superimposed voids (grey regions). A new tendency (red vector) is computed as an average of reference tendencies (black vectors) in the neighborhood of sampled SST (red contour).
In order to sample from the JPD we use the inverse transform sampling method (Devroye, 1986). We have also tried the rejection sampling but did not gain that much of a difference in its favour; detailed explanations on sampling from the JPD are given in (Shevchenko and Berloff, 2023b). Sampling from the JPD allows to probabilistically reconstruct the originally sparsely-populated reference phase space by filling it with new locations and directional vectors (tendencies) needed to advect the image point (Fig. 1d). Note that the probabilistically-reconstructed reference data can be used not only in the HP method, but in other data-driven methods which suffer lack of data.
It is important to note that SST and its tendencies in the probabilistically-reconstructed phase space are not ordered in time. This is the moment when we need Advection of the image point to compute a trajectory (solution), i.e to order those points in time.
(e) Advection of the image point in probabilistically-reconstructed phase space. In the fifth step, the probabilistically-reconstructed phase space is used by Advection of the image point to calculate the trajectory (red line in Fig. 1e). For the reader’s convenience, we briefly describe how Advection of the image point works. This method falls into the category of data-driven methods from the hyper-parametersation (HP) class which takes advantage of working in phase spaces, as opposed to the conventional methods operating in the physical space. The method has been tested on both idealized and comprehensive ocean models (two-layer quasi-geostrophic model in a channel and MITgcm in the North Atlantic configuration) and showed significant improvements of the HP solution toward the reference one (Shevchenko and Berloff, 2021, 2023a). The HP approach currently umbrellas five different methods (ranging from purely data-driven to hybrids, which combine the data- and physics-driven paradigms). Another striking feature of the HP approach is that its measure of goodness is how close the HP solution is to the reference phase space. To put it another way, the HP approach matches phase spaces (the reference phase space and the phase space where the HP solution evolves), whereas conventional methods match individual trajectories. This measure allows the HP solution to evolve in the neighborhood of the reference phase space, and therefore reproduce the flow dynamics which is very similar to the reference one.
The idea behind Advection of the image point (which we refer to as the HP method in what follows) is to use reference data, say $\mathbf{x}\in\mathbb{R}^{n}$, and the following differential equation
$\frac{d\mathbf{y}}{dt}=\frac{1}{M}\sum\limits_{j=\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})\Big{|}_{\mathcal{U}(\mathbf{y}(t))}+\eta\left(\frac{1}{N}\sum\limits_{i=\mathcal{U}_{I}}\mathbf{x}_{i}\Big{|}_{\mathcal{U}(\mathbf{y}(t))}-\mathbf{y}(t)\right),\quad\mathbf{y}(t_{0})=\mathbf{y}_{0},$ (1)
to describe the evolution of the image point, $\mathbf{y}\in\mathbb{R}^{n}$, in the phase space of reference data. In the context of this work, $\mathbf{x}(t)$ is the probabilistically-reconstructed SST, and $\mathbf{y}(t)$ is the HP solution computed in the probabilistically-reconstructed phase space, and $\eta$ is the nudging strength. The neighbourhood of $\mathbf{y}(t)$ is denoted as $\mathcal{U}(\mathbf{y}(t))$, $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ are the sets of timesteps indexing the discrete reference solution $\mathbf{x}$ in $\mathcal{U}(\mathbf{y}(t))$. Although the method allows to use different sets of reference tendencies $\mathbf{F}(\mathbf{x}_{j})$ and points $\mathbf{x}_{i}$, we set $M=N$ and $\mathcal{U}_{I}=\mathcal{U}_{J}$ in this study.
In a nutshell, the actual dynamics is given by the observed reference tendency $\frac{1}{M}\sum\nolimits_{j=\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})\Big{|}_{\mathcal{U}(\mathbf{y}(t))}$ the imperfect reconstruction of which from the reference data is compensated by nudging towards the observed reference state $\frac{1}{N}\sum\nolimits_{i=\mathcal{U}_{I}}\mathbf{x}_{i}\Big{|}_{\mathcal{U}(\mathbf{y}(t))}$ in the probabilistically-reconstructed phase space. Note that the neighbourhood is computed as the average over $M$ (and $N$ for the nudging term) nearest, in $l_{2}$ norm, to the solution $\mathbf{y}(t)$ points. The choice of the norm and the way the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is calculated is not limited to those used in this study, and can vary depending on what is needed. Our choice is, probably, the simplest, but it serves well the purpose of this work. We dub the set of parameters $\{M,N,\eta\}$ hyper-parameters. The hyper-parameters can be set based on the chosen metric and available data, we will return to their choice later. The interested reader is referred to (Shevchenko and Berloff, 2021, 2023a) for a more detailed discussion of the method.
Figure 3: Shown is a sketch of the probabilistically-reconstructed reference phase space (orange blob), vectors $\mathbf{F}(\mathbf{x}_{i})$ (black arrows) pointing out from points $\mathbf{x}_{i}$. For the HP method all these points is a cloud of points which are not ordered in time. The HP solution $\mathbf{y}(t)$ is shown as a red trajectory, and the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is the black contour surrounding the red dot, which is the HP solution $\mathbf{y}(t)$ at time $t$. The red vectors (and also every point they are attached to) within the neighborhood are the $N$ reference fields $\mathbf{x}_{i},\,i\in\mathcal{U}_{I}$ (and $M$ reference tendencies $\mathbf{F}(\mathbf{x}_{i})$) nearest, in $l_{2}$-norm, to the HP solution $\mathbf{y}(t)$. We assume that $\mathcal{U}_{I}=\mathcal{U}_{J}$ and $M=N$, but the method is not bound to this particular choice.
The characteristic feature of the HP method is that it looks at the reference dataset as being a cloud of points which are not ordered in time. This is why we write $\mathbf{x}_{i}$ and $\mathbf{F}(\mathbf{x}_{j})$ instead of $\mathbf{x}(t)$ and $\mathbf{F}(\mathbf{x}(t))$ in equation (1). At every time step, the HP method takes $N$ nearest points $\mathbf{x}_{i}$ (and $M$ tendencies $\mathbf{F}(\mathbf{x}_{j})$) to the solution $\mathbf{y}(t)$ points (Fig. 3). Thus, the HP method never runs out of reference data, as a set of nearest points to the HP solution $\mathbf{y}(t)$ can always be found. The main advantage of using the HP method in probabilistically-reconstructed phase spaces is that the data for the method is sampled from the same distribution as the reference data itself.
Calibration of hyper-parameters. The HP method has hyper-parameters $\{M,N,\eta\}$ which should be properly set to address the problem at hand. In order to calibrate these hyper-parameters we define a measure of discrepancy of two clouds of points, say $A$ and $B$, as follows
$\mathrm{Di}(A,B):=\|A-\{\widehat{\min}\,d(a_{i},B)\}_{1\leq i\leq|A|}\|_{2}/\|A\|_{2}\,,$ (2)
where $d$ is the Euclidean distance between the $i$-th point in $A$ and a point in $B$; ||$\cdot$|| ⋅ | is the number of points in the cloud (it is assumed that $|A|=|B|$). The cloud A represents the reference data, while the cloud B is the HP solution computed in the probabilistically-reconstructed phase space; the clouds are not supposed to have any order defined on them.
It is important to remark that the $\widehat{\min}$ in (2) is excluding, i.e. once the minimum between a pair $(a_{i},b_{j})$ has been found, the point $b_{j}$ is excluded from $B$ (to avoid multiple counting) and is then used in comparison with element $a_{i}$. In other words, $\widehat{\min}$ does not return the minimum distance between $a_{i}$ and $B$ per se, it returns the element $b_{j}$ that delivers this minimum distance. Also note that the choice of this measure of discrepancy reflects the measure of goodness used for the HP method (namely, the proximity of the phase space (where HP solution evolves) to the reference phase space, or more accurately to the probabilistically-reconstructed reference phase space). The calibration procedure consists of minimizing the discrepancy on the space of hyper-parameters; the smaller the discrepancy is, the closer the clouds are to each other. We would like to remind the reader again that the HP method tries to match phase spaces (not individual trajectories), and this is why the calibration of hyper-parameters makes itself useful.
(f) Evolution in physical space. The fifth step is to get the modelled solution back in the physical space (Fig. 1f). In this study, the solution computed with the HP method is reshaped from a vector (used in the phase space) into a matrix, which is used to present fields in the physical space.
3 The probabilistic reconstruction method in action
In this section we apply the proposed method to Chua’s circuit to demonstrate how it works in a low-dimensional phase space and then consider its application in the context of the global NEMO model.
3.1 Chua’s circuit
As a minimum working example, we consider Chua’s circuit and demonstrate in more detail how the method works. Chua’s circuit (Chua et al., 1986) is given by the following system of equations:
$\frac{d\mathbf{x}}{dt}=\mathbf{F}(\mathbf{x}(t)),\quad\mathbf{F}:=\begin{pmatrix}\alpha(y-ax^{3}-cx)\\ x-y+z\\ -\beta y-\gamma z\\ \end{pmatrix},$ (3)
with $\mathbf{x}(t)=(x(t),y(t),z(t))$, and $\alpha=10$, $\beta=15$, $\gamma=0.01$, $a=0.1$, $c=-0.2$. As an initial condition, we take $\mathbf{x}(t_{0})=(1,0,0)$ so that the solution is close to the attractor.
Reference data acquisition. Within the context of Chua’s circuit, the reference data for the HP method is the solution $\mathbf{x}(t)$ and its tendencies $\mathbf{F}(\mathbf{x}(t))$. In order to get the reference data we integrate the Chua system (3) over time $t\in[0,100]$ and save the solution. Then, the tendencies are computed by differentiating $\mathbf{x}(t)$; we use the central finite difference in time (Fig. 4). Instead of differentiating the solution, one can use the right hand side, but we recommend against doing that, as the right hand side is usually unavailable for numerical simulations and observations. Note that instead of or in combination with the numerical solution, one can also use observations. In this case, one should form a new dataset $\widehat{\mathbf{x}}=\{\mathbf{x}(t),\mathbf{x}_{o}(t)\}$, which includes both the solution $\mathbf{x}(t)$ and observations $\mathbf{x}_{o}(t)$.
Figure 4: Shown is (a) the solution of Chua’s system $\mathbf{x}(t)$ and (b) the corresponding vector field $\mathbf{F}(\mathbf{x}(t))$ for $t=[0,100]$.
Calculation of JPD and sampling from JPD. The JPD of the reference solution $\mathbf{x}(t)$ and reference tendencies $\mathbf{F}(\mathbf{x}(t))$ are calculated as described above. In order to make sure that sampling from the JPD is accurate enough for further use, we compare the reference PDF (computed from the reference dataset) and the reconstructed PDF (computed from the data sampled from the JPD). As seen in Fig. 5, the reconstructed PDF (blue) is an accurate approximation to the reference PDF (black); the reconstructed PDF can be calculated more accurately, but it is not required for the purpose of this study.
Figure 5: Shown is the coordinate-wise reference PDF (black) and reconstructed PDF (blue) for the solution of Chua’s system (top row) and for the corresponding vector field (bottom row). Note that the reconstructed PDF is an accurate approximation to the reference PDF.
The next step is to sample from the JPD to probabilistically reconstruct the reference phase space for the HP method. We use the same amount of reference data as before, but harvest twice as much from the JPD (Figs. 6,7). The results clearly demonstrate the high quality of the probabilistically-reconstructed data, and that it covers a larger phase space compared with the reference one.
Figure 6: Shown are different projections of the reference solution (top row) and its reconstruction from the JPD (bottom row). The reconstructed data not only reproduces the reference one but, more importantly, densifies the phase space by adding new points, as we sampled twice as much points from the JPD compared with the reference data.
Figure 7: The same as in Fig. 6, but for the tendencies. Note that the tendencies are not sampled directly from the JPD but are computed locally in the neighbourhood of the sampled solution.
3.2 Calibration of hyper-parameters
The calibration procedure consists of minimizing $\mathrm{Di}(A,B)$ in the space of hyper-parameters $M$, $N$, and $\eta$. One can opt for engaging a full-blown optimization to find the optimum solution. However, for the purpose of this study, we set $M=N=5$ and search for $\eta\in[0,1]$ that minimizes $\mathrm{Di}(A,B)$. We have found that $\eta=0.235$ gives a solution with low discrepancy.
In order to demonstrate that the proposed method can gain from using the probabilistically-reconstructed phase space, we compare it with the HP method run on the reference dataset (Fig. 8).
Figure 8: Shown are (a) the reference solution, (b) HP solution computed in the probabilistically-reconstructed reference phase space, and (c) HP solution computed in the reference phase space (no probabilistic reconstruction used); both HP solutions are computed for $M=N=5$ and $\eta=0.235$. The HP solution computed in the probabilistically-reconstructed reference phase space is much more developed compared with the HP solution computed in the reference phase space. Note that both HP solutions run for $t\in[0,200]$, i.e. 50% of the time out of the reference sample.
As seen in Fig. 8b, the HP solution computed in the probabilistically-reconstructed phase space is very-well developed on the attractor compared with the HP solution computed on the reference dataset (Fig. 8c), which remains bounded to a very narrow band because of lack of data. All these results give an extra layer of confidence that the proposed method has strong potential for the use in more sophisticated models which we consider in the next section.
3.3 NEMO model
The output of the global 1/4 (ORCA025-N4001) and 1/12 (ORCA0083-N006) resolution NEMO model is used in this study. The reference solution is the 5-day mean sea surface temperature (SST) and surface relative vorticity (SRV) in the North Atlantic region $[-83^{\circ}W,-20^{\circ}W]\times[27^{\circ}N,55^{\circ}N]$ over the period of 1979-1993; the period of 1994-2007 is used for validation to check how the HP method performs. The reference solution is interpolated from the 1/12-grid onto the 1/4 grid. The lower-resolution 1/4-solution from ORCA025-N4001 simulation (we call it the modelled solution) is used for comparison with the HP solution.
In order to calibrate the hyper-parameters we take $M=N=5$ and minimize $\mathrm{Di}(A,B)$ with respect to $\eta$. We have found that $\eta=0.01$ gives a low discrepancy. After setting up the hyper-parameters, we compute a probabilistically-reconstructed phase space by sampling twice as much compared to the reference record length (Fig. 9).
Figure 9: Shown are typical SST (top row) and SRV (bottom row) fields sampled from the reference JPD.
The dimensionality of the probabilistically-reconstructed space is $n\approx 30000$. One can run the HP method in the high-dimensional phase space, but in this work we further reduce it down to $m=200$ by using the EOF-PC decomposition (Preisendorfer, 1988). It results in a two-order of magnitude reduction compared with the original dimension; note that the 200 leading EOFs capture 99% of the total SST variance and 60% of the total SRV variance. We compute the EOF-PC decomposition of SST and SRV separately. However, it can be computed for both if one needs to model these fields together.
The space of PCs (let it be denoted as $\mathbf{x}$) can be regarded as a reduced version of the probabilistically-reconstructed phase space. Given this reduced space, one can reconstruct a dynamical system governing the reduced dynamics of observed reference data, as we did in (Shevchenko and Berloff, 2022b). In this study, we use a different idea: the dynamics of PCs is modelled in the reduced space with the HP method directly and then the HP solution is projected back into the full-dimensional probabilistically-reconstructed phase space.
In the reduced phase space, equation (1) reads as
$\frac{d\mathbf{z}}{dt}=\frac{1}{M}\sum\limits_{j=\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})\Big{|}_{\mathcal{U}(\mathbf{z}(t))}+\eta\left(\frac{1}{N}\sum\limits_{i=\mathcal{U}_{I}}\mathbf{x}_{i}\Big{|}_{\mathcal{U}(\mathbf{z}(t))}-\mathbf{z}(t)\right),\quad\mathbf{z}(t_{0})=\mathbf{z}_{0}.$ (4)
The only difference of (4) compared with equation (1) is that $\mathbf{x}$ does not represent the reference solution (it would be SST or SRV in this case) but the PCs. Having solved equation (4), we approximate the HP solution, $\mathbf{y}(t)$, in the probabilistically-reconstructed phase space by using the leading EOF-PC pairs as follows:
$\mathbf{y}(t)\approx\sum\limits^{m}_{i=1}z_{i}(t)\mathbf{E}_{i}\,,$ (5)
with $\mathbf{E}_{i}$ and $z_{i}$ being the i-th EOF and PC (computed with the HP method in the reduced space), respectively. Thus, solution $\mathbf{y}(t)$ is the SST (or SRV) computed in the full-dimensional probabilistically-reconstructed phase space.
We report the results in Fig. 10 which clearly shows that the HP solution (Fig. 10b) much better represents the Gulf Stream than the modelled solution (Fig. 10c). These results are also confirmed in Fig. 11. Namely, the difference between the time mean of the reference solution and HP solution is much smaller (Fig. 11a) than that of the modelled solution (Fig. 11b). The root mean square error for the HP solution (Fig. 11c) is also much lower then the one of the modelled solution (Fig. 11d).
Figure 10: Shown are typical SST snapshots at 1/4 resolution for (a) the reference solution, (b) HP solution, and (c) modelled solution. The HP solution significantly better represents the Gulf Stream compared with the modelled solution.
Figure 11: Shown are (a) difference between the time mean reference SST and HP solution, (b) difference between the time mean reference and modelled SST, (c) root mean square error between the reference SST and HP solution, (d) root mean square error between the reference and modelled SST. We use $m=200$ in the EOF-PC decomposition.
More insights about how the HP solution differs from the modelled solution can be seen in the SRV fields (Fig. 12). The Gulf Stream computed with the HP method (Fig. 12b) is much stronger than the one computed with the 1/4 model (Fig. 12c). Moreover, the HP solution is teemed with vortices (like the reference one), while the vortex dynamics is significantly inhibited in the modelled solution as it is also reflected by the error plots in Fig. 13. Doubling the number of PCs ($m=400$) results in a more pronounced Gulf Stream dynamics and larger population of vortices (Fig. 12). However it is contribution to the time mean flow and the root mean square error is small (not shown).
Figure 12: The same as in Fig. 11 but for SRV ($m=200$ and $m=400$).
Figure 13: The same as in Fig. 11 but for SRV.
4 Conclusions
In this study we have proposed to use the joint probability distribution (JPD) to probabilistically reconstruct the phase space which can be used when the original reference phase space (computed from numerical solutions, observations, or both) is too sparse to allow the use of data-driven methods. The probabilistically-reconstructed phase space is calculated by sampling from the JPD, thus providing an extra source of data from the reference distribution and ensuring that both the reference phase space and the reconstructed one share the same reference distribution. We have shown how the probabilistic reconstruction works on the example of Chua’s system and then applied it within the context of the global ocean model NEMO. We have also found that the HP method can be used directly in reduced phase spaces thus allowing several-orders-of-magnitude acceleration compared to the 1/4-resolution NEMO run. We reduced the dimensionality of the reference phase space by two orders of magnitude by applying the EOF-PC decomposition.
Our results show that the use of the HP method in the probabilistically-reconstructed (and reduced) phase space gives more accurate results and more realistic flow dynamics compared with the 1/4-NEMO simulation. It is also worth mentioning that the two order of magnitude acceleration can be translated into using larger ensemble of HP solutions for probabilistic predictions.
We believe that the proposed method demonstrates the strong potential for the use in operational ocean and ocean-atmospheric models. The next step in this direction is to assess prediction skills and achievable forecast range of the proposed method for high-resolution forecasting systems and near-term climate predictions.
Acknowledgements.
[…]